# Group Therapy for Halos: Advancing Halo Mass Estimation for Galaxy Groups

### v1

@@ Abstract
Accurate estimation of dark matter halo masses for galaxy groups is central to studies of galaxy evolution and for leveraging group catalogues as cosmological probes. In this work, we present a comprehensive calibration and evaluation of two complementary halo mass estimators: a dynamical estimator based on the virial theorem, and an empirical relation between the sum of the stellar masses of the three most massive group galaxies and the halo mass (SHMR). Using a suite of state-of-the-art semi-analytic models (SAMs; Shark, SAGE, and GAEA) to produce observationally motivated mock light-cone catalogues, we rigorously quantify the accuracy, uncertainty, and model dependence of each method. The calibrated virial theorem achieves negligible systematic bias (mean $\Delta=-0.01$ dex) and low scatter (mean $\sigma=0.20$ dex) as a function of the predicted halo mass, with no sensitivity to the SAM baryonic physics. The calibrated SHMR yields the highest precision, with mean $\Delta=0.02$ dex and mean $\sigma=0.14$ dex as a function of the predicted halo mass but exhibits greater model dependence due to its sensitivity to varying baryonic physics and physical prescriptions across the SAMs. We demonstrate the application of these estimators to observational group catalogues, including the construction of the empirical halo mass function and the mapping of quenched fractions in the stellar mass–halo mass plane. We provide clear guidance on the optimal application of each method: the calibrated virial theorem is recommended for GAMA-like surveys ($i<19.2$) calibrated to $z<0.1$ and should be used for studies requiring minimal model dependence, while the SHMR is optimal for high-precision halo mass estimation across diverse catalogues with magnitude limits of $Z<21.2$ or brighter and to redshifts of $z\leq 0.3$. These calibrated estimators will be of particular value for upcoming wide-area spectroscopic surveys, enabling robust and precise analyses of the galaxy–halo connection and the dark matter halo population.
@@ 1 Introduction
The dark-matter halo-mass represents one of the most fundamental properties governing galaxy evolution within the hierarchical structure formation paradigm; as galaxies form and evolve within these gravitationally bound dark matter structures, the halo mass strongly influences numerous galaxy properties including star formation histories, morphologies, and chemical enrichment pathways (White & Rees, 1978; Blumenthal et al., 1984; Bower et al., 2006). Accurate determination of halo masses therefore underpins our ability to connect theoretical models of structure formation with observational constraints on galaxy evolution across cosmic time.
Given the importance of halo mass in the field of galaxy evolution, accurate measures of halo mass are critical; however, it remains challenging to measure directly. Unlike baryonic matter, dark matter does not emit electromagnetic radiation, leaving astronomers to rely on indirect probes such as gravitational lensing (e.g. Mandelbaum et al., 2016), X-ray emission from hot gas (e.g. Arnaud et al., 2010), or dynamical tracers like galaxy kinematics (e.g. Old et al., 2014). Each method is subject to significant limitations and systematic uncertainties, which are particularly acute at the group scale where the number of galaxy members is low. This regime is crucial for tracing the growth of structure within group environments and for understanding the transition from group to cluster environments (Yang et al., 2007; Robotham et al., 2011).
@@
Current approaches to estimating halo masses face several challenges: weak lensing signals become increasingly difficult to detect for lower-mass systems (Viola et al., 2015), while X-ray observations require the presence of hot gas in hydrostatic equilibrium—a condition typically not met in group-scale environments (Lovisari et al., 2021). Traditional dynamical mass estimators based on the virial theorem assume that systems are both virialised and well-sampled; these assumptions often break down for groups with low multiplicity, particularly when survey completeness is limited (Robotham et al., 2011; Old et al., 2018; Wojtak et al., 2018). Collectively, these limitations introduce substantial scatter and systematic uncertainty in halo mass estimates, particularly in the low mass/multiplicity group regime where accurate halo mass measurements are most critical for understanding the impact of baryonic feedback on galaxy evolution (Wechsler & Tinker, 2018).
@@
Improved halo mass estimations are required to address several fundamental questions in modern astrophysics. First, the precise shape and amplitude of the halo mass function provides important constraints on cosmological parameters, particularly $\sigma_{8}$ and $\Omega_{m}$ (Tinker et al., 2008; Castro et al., 2021; Driver et al., 2022a). Second, understanding the coevolution of galaxies and their host haloes requires accurate mapping between observable galaxy properties and underlying halo masses (Behroozi et al., 2019; Moster et al., 2020). The scatter in the stellar halo mass relation (SHMR) offers insights into the complex baryonic processes driving galaxy assembly, including feedback from supernovae, active galactic nuclei (AGN), and star-formation efficiency (Pillepich et al., 2018; Davies et al., 2019; Oyarzún et al., 2024; Wang & Peng, 2025). More broadly, measuring the baryon content of haloes as a function of halo mass remains a central goal in galaxy evolution and cosmology. For example, Chauhan et al. (2020) demonstrated that the signatures of AGN feedback on the HI content of haloes are remarkably strong, with variations between simulations exceeding 1 dex. However, subsequent work by Chauhan et al. (2021) showed that the uncertainties inherent in halo mass estimation, particularly when using HI stacking techniques commonly adopted in the literature, the uncertainties in halo mass can completely obscure these feedback signatures, making it impossible to distinguish between different simulation models. These studies further highlight that while dynamical mass estimators perform well for high-multiplicity systems, conversions based on stellar mass or luminosity provide more reliable halo mass estimates for low-multiplicity groups. This underscores the critical importance of minimising halo mass uncertainties in order to robustly connect baryonic content and feedback processes to the underlying dark matter halo population.
@@
The advent of next-generation wide-field spectroscopic surveys such as Wide Area Vista Extragalactic Survey (WAVES; Driver et al., 2019), DESI (DESI Collaboration et al., 2016), and the 4MOST Hemisphere Survey (Taylor et al., 2023) promises to generate comprehensive catalogues of galaxy groups across cosmic time. These surveys will deliver spectroscopic redshifts for millions of galaxies, where high completeness will facilitate the robust identification of gravitationally bound structures at the group scale where environmental effects demonstrably alter galaxy properties and evolutionary trajectories (Peng et al., 2010; Wetzel et al., 2013; Davies et al., 2019). However, the scientific potential of these datasets critically depends on our capacity to accurately translate observable group properties into reliable halo mass estimates (Kravtsov et al., 2018; Eckert et al., 2020; Tinker, 2021). This connection between observable baryonic tracers and the underlying dark matter distribution remains fundamental for quantitatively testing hierarchical structure formation models.
@@
The continuous flow of gas into, within, and out of galaxies is referred to as the baryon cycle and represents a key process governing galaxy evolution (Davé et al., 2012; Lilly et al., 2013). Complementary multi-wavelength facilities such as Euclid (Laureijs et al., 2011), the Vera C. Rubin Observatory (LSST Science Collaboration et al., 2009), the Square Kilometre Array (Dewdney et al., 2009), ALMA (Wootten & Thompson, 2009), eROSITA (Merloni et al., 2012), ASKAP (Johnston et al., 2008), and VLT/MUSE (Bacon et al., 2010) will capture distinct components of this cycle: stellar content through optical/near-infrared observations, molecular gas reservoirs via millimeter observations, hot gas through X-ray measurements, neutral hydrogen via radio observations, and spatially resolved gas kinematics through integral field spectroscopy (Saintonge et al., 2017; Péroux & Howk, 2020; Tacconi et al., 2020). Pairing these multi-wavelength datasets with spectroscopic information and group catalogues allows us to directly probe the influence of dark matter on the baryon cycle and empirically test how halo properties regulate key processes such as gas accretion rates, star formation efficiency, and feedback-driven outflows across diverse environments and cosmic epochs (Tumlinson et al., 2017; Mitchell et al., 2020; van de Voort et al., 2021). A comprehensive understanding of these complex relationships necessitates both the statistical power of large-scale spectroscopic group catalogues and the detailed multi-wavelength characterisation of baryon cycle components. Nevertheless, precise halo mass measurements remain the critical prerequisite, particularly at the group scale ($10^{12}$$10^{14}$ M) where the interplay between dark matter and baryonic processes most significantly influences galaxy evolution (Behroozi et al., 2019; Davies et al., 2019; Wang & Peng, 2025).
@@
Semi-analytical models (SAMs) offer a powerful framework for developing and calibrating such techniques. These models implement physically motivated prescriptions for galaxy formation within dark matter halo merger trees from N-body simulations (Prada et al., 2012; Somerville et al., 2015; Croton et al., 2016; Lagos et al., 2018; De Lucia et al., 2024). By producing realistic galaxy populations with known halo properties, SAMs allow us to assess the performance of different mass estimation techniques and quantify their associated uncertainties. Recent advances in SAMs, including improved treatments of gas cooling, star formation, and feedback processes, which has significantly enhanced their ability to reproduce observed galaxy properties across a wide range of environments (Klypin et al., 2016; Croton et al., 2016; Lacey et al., 2016; Lagos et al., 2018; Stevens et al., 2018; Henriques et al., 2020; Lagos et al., 2024; De Lucia et al., 2024).
@@
In this paper, we present two complementary approaches to improve observational halo mass estimates for galaxy groups. The first approach addresses limitations in traditional virial mass estimators when applied to low-multiplicity groups by developing a corrective framework that accounts for systematic biases in groups with small velocity dispersion and projected radius measurements. The second method utilises the correlation between baryonic mass and dark matter mass, probing the mass relationship between the 3 most massive galaxies in the halo and that of the halo mass.
@@
This paper is organised as follows. In Section 2, we describe the semi-analytical models used to calibrate/compare our relations and introduce the observational dataset in which we showcase applications of the relations. Section 3 introduces our two calibrated halo mass estimators; Section 3.1 presents the first approach utilising a modified virial theorem to estimate halo mass, and in Section 3.2 we present a calibrated SHMR using the 3 most massive galaxies in a halo as a proxy for halo mass. In Section 4, we apply the halo mass estimations to the observational data and demonstrate their performance and use cases. Finally, Section 5 summarises our findings and outlines future directions for halo mass estimation in upcoming large-scale surveys.
@@
2 Data and Simulations
This section provides a comprehensive overview of the simulated and observational datasets utilised in this study. We first describe the suite of state-of-the-art semi-analytic models (SAMs) of galaxy formation employed for calibration and validation, including Shark, SAGE, and GAEA, in Sections 2.1, 2.1.2 and 2.1.3 respectively, each of which incorporates distinct treatments of baryonic physics and galaxy evolution processes. We then introduce the Southern Galactic Pole (SGP) dataset and its group catalogue, which serves as the primary observational dataset for the application and testing of the calibrated halo mass estimators (Section 2.2).
2.1 Semi-Analytic Models of Galaxy Formation
@@
SAMs represent powerful theoretical tools for exploring the complex interplay between dark matter structure formation and baryonic physics. These models implement physically motivated prescriptions for key astrophysical processes within merger trees extracted from cosmological N-body simulations (see reviews by: Baugh, 2006; Benson, 2010). By providing self-consistent galaxy populations with fully known dark matter halo properties, SAMs offer an ideal framework for developing and calibrating halo mass estimation techniques. In this work, we employ three independent SAMs: Shark (Lagos et al., 2018), SAGE (Croton et al., 2016) and GAEA (De Lucia et al., 2024), which implement different physical prescriptions while operating on distinct N-body simulations. This multi-model approach allows us to assess the robustness of our halo mass estimation methods against variations in the underlying galaxy formation physics.
@@
To ensure that the simulated data are directly comparable to observations, each SAM output was processed with an explicitly observational mindset. Mock light cones were constructed by projecting the simulated volumes into realistic sky coordinates, redshifts, and cosmological distances. By incorporating galaxy peculiar velocities, the resulting mock redshifts naturally include redshift-space distortions, such as the finger-of-god and Kaiser effects, thereby replicating key observational systematics. Galaxy group samples were intrinsically constructed by matching galaxies to their host halo IDs, allowing us to assess the performance of halo mass estimators in “perfect” groups whilst still subjecting them to observational biases. Apparent magnitudes were computed for each galaxy, and a Galaxy and Mass Assembley (GAMA) G23 magnitude limit of $i<19.2$ was imposed on the mock light cones to provide observational-like magnitude limits (Driver et al., 2009, 2011). This methodology ensures that the calibration and validation of halo mass estimators are performed under conditions that closely mimic those encountered in real, highly complete spectroscopic surveys, thereby enhancing the reliability and applicability of the methods to current and future observational datasets. As a caveat, a fainter magnitude limit of $Z<21.2$ was utilised for SHARK in the development of the SHMR (see Section 3.2.3 for more details).
2.1.1 SHARK

### v3

@@ Abstract
Accurate estimation of dark matter halo masses for galaxy groups is central to studies of galaxy evolution and for leveraging group catalogues as cosmological probes. In this work, we present a comprehensive evaluation and calibration of two complementary halo mass estimators: a dynamical estimator based on a modified virial theorem (MVT), and an empirical summed stellar mass to halo mass relation (sSHMR) which uses the summed mass of the three most massive group galaxies as a proxy for halo mass. Using a suite of state-of-the-art semi-analytic models (SAMs; Shark, SAGE, and GAEA) to produce observationally motivated mock light-cone catalogues, we rigorously quantify the accuracy, uncertainty, and model dependence of each method. The MVT halo mass estimator achieves negligible systematic bias (mean $\Delta=-0.01$ dex) and low scatter (mean $\sigma=0.20$ dex) as a function of the predicted halo mass, with no sensitivity to the SAM baryonic physics. The calibrated sSHMR yields the highest precision, with mean $\Delta=0.02$ dex and mean $\sigma=0.14$ dex as a function of the predicted halo mass but exhibits greater model dependence due to its sensitivity to varying baryonic physics and physical prescriptions across the SAMs. We demonstrate the application of these estimators to observational group catalogues, including the construction of the empirical halo mass function and the mapping of quenched fractions in the stellar mass–halo mass plane. We provide clear guidance on the optimal application of each method: the MVT is recommended for GAMA-like surveys ($i<19.2$) calibrated to $z<0.1$ and should be used for studies that require minimal model dependence, while the sSHMR is optimal for high-precision halo mass estimation across diverse catalogues with magnitude limits of $Z<21.2$ or brighter and to redshifts of $z\leq 0.3$. These calibrated estimators will be of particular value for upcoming wide-area spectroscopic surveys, enabling robust and precise analyses between the galaxy–halo connection and the underlying dark matter distribution.
@@ 1 Introduction
The dark-matter halo-mass represents one of the most fundamental properties governing galaxy evolution within the hierarchical structure formation paradigm; as galaxies form and evolve within these gravitationally bound dark matter structures, the halo mass strongly influences numerous galaxy properties including star formation histories, morphologies, and chemical enrichment pathways (White1978; Blumenthal1984; Bower2006). Accurate determination of halo masses therefore underpins our ability to connect theoretical models of structure formation with observational constraints on galaxy evolution across cosmic time.
@@
Given the importance of halo mass in the field of galaxy evolution, accurate measures of halo mass are critical; however, it remains challenging to measure directly for individual groups. Weak gravitational lensing provides accurate and widely adopted halo mass estimates, and is routinely used to validate indirect estimators (Leauthaud2012; Hoekstra2013; Velander2014; Mandelbaum2016; Mandelbaum2018). Nonetheless, robust weak-lensing halo masses for individual galaxy groups are not universally available at typical survey depths, as the per-object signal-to-noise at group-scale masses is often insufficient without stacking or targeted deep imaging (Leauthaud2012; Velander2014; Viola2015; Simet2017). Other indirect probes include X-ray emission from hot gas (e.g., Arnaud2010) and galaxy dynamics (e.g., Old2014); these can be applied to individual groups, but carry method-specific systematics that are particularly pronounced in low membership groups. This regime is crucial for tracing the growth of structure within group environments and for understanding the transition from group to cluster environments (Yang2007; Robotham2011). Current approaches to estimating halo masses face several challenges: weak lensing signals become increasingly difficult to detect for lower-mass systems (Viola2015), while X-ray observations require the presence of hot gas in hydrostatic equilibrium; a condition typically not met in group-scale environments (Lovisari2021). Traditional dynamical mass estimators based on the virial theorem assume that systems are both virialised and well-sampled; these assumptions often break down for groups with low multiplicity, particularly when survey completeness is limited (Robotham2011; Old2018; Wojtak2018). Collectively, these limitations introduce substantial scatter and systematic uncertainty in halo mass estimates, particularly in the low mass/multiplicity group regime where accurate halo mass measurements are most critical for understanding the impact of baryonic feedback on galaxy evolution (Wechsler2018).
@@
Improved halo mass estimations are required to address several fundamental questions in modern astrophysics. First, the precise shape and amplitude of the halo mass function provides important constraints on cosmological parameters, particularly $\sigma_{8}$ and $\Omega_{m}$ (Tinker2008; Castro2021; Driver2022). Second, understanding the coevolution of galaxies and their host haloes requires accurate mapping between observable galaxy properties and underlying halo masses (Behroozi2019; Moster2020). The scatter in the stellar halo mass relation (SHMR) offers insights into the complex baryonic processes driving galaxy assembly, including feedback from supernovae, active galactic nuclei (AGN), and star-formation efficiency (Pillepich2018; Davies2019; Oyarzun2024; Wang2025). More broadly, measuring the baryonic properties of galaxies as a function of halo mass remains a central goal in galaxy evolution and cosmology. For example, Chauhan2020 demonstrated that the signatures of AGN feedback on the HI content of haloes are remarkably strong, with variations between simulations exceeding 1 dex. However, subsequent work by Chauhan2021 showed that the uncertainties inherent in halo mass estimation, particularly when using the HI stacking techniques commonly adopted in the literature, can completely obscure these feedback signatures, makes it impossible to distinguish between different simulation models. These studies further highlight that while dynamical mass estimators perform well for high-multiplicity systems, but are unable to resolve the underlying distribution for low mass, low multiplicity groups. This underscores the critical importance of minimising halo mass uncertainties in order to robustly connect baryonic content and feedback processes to the underlying dark matter halo population.
@@
The advent of next-generation wide-field spectroscopic surveys such as Wide Area Vista Extragalactic Survey (WAVES; Driver2019), the Dark Energy Spectroscopic Instrument (Desi; DESI2016), and the 4MOST Hemisphere Survey (ENTaylor23) promises to generate comprehensive catalogues of galaxy groups across cosmic time. These surveys will deliver spectroscopic redshifts for millions of galaxies, where high completeness will facilitate the robust identification of gravitationally bound structures at the group scale where environmental effects demonstrably alter galaxy properties and evolutionary trajectories (Peng2010; Wetzel2013; Davies2019). However, the scientific potential of these datasets precariously depends on our capacity to accurately translate observable group properties into reliable halo mass estimates (Kravtsov2018; Eckert2020; Tinker2021). This connection between observable baryonic tracers and the underlying dark matter distribution remains fundamental for quantitatively testing hierarchical structure formation models.
@@
The continuous flow of gas into, within, and out of galaxies is referred to as the baryon cycle and represents a key process governing galaxy evolution (Davé2012; Lilly2013). Complementary multi-wavelength facilities such as Euclid (Laureijs2011), the Vera Rubin Observatory (LSST2009), the Square Kilometre Array (SKA; Dewdney2009), the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten2009), the extended Roentgen Survey with an Imaging Telescope Array (eROSITA; Merloni2012), the Australian Square Kilometre Array Pathfinder (ASKAP; Johnston2008), and the Very Large Telescope/Multi Unit Spectroscopic Explorer (VLT/MUSE; Bacon2010) capture distinct components of this cycle: stellar content through optical and near-infrared observations, molecular gas reservoirs via millimetre observations, hot gas through X-ray measurements, neutral hydrogen via radio observations, and spatially resolved gas kinematics through integral-field spectroscopy (Saintonge2017; Peroux2020; Tacconi2020). Pairing these multi-wavelength datasets with spectroscopic information and group catalogues allows us to directly probe the influence of dark matter on the baryon cycle and empirically test how halo properties regulate key processes such as gas accretion rates, star formation efficiency, and feedback-driven outflows across diverse environments and cosmic epochs (Tumlinson2017; Mitchell2020; van_de_Voort2021). A comprehensive understanding of these complex relationships necessitates both the statistical power of large-scale spectroscopic group catalogues and the detailed multi-wavelength characterisation of baryon cycle components. Nevertheless, precise halo mass measurements remain the critical prerequisite, particularly at the group scale ($10^{12}$$10^{14}$ M) where the interplay between dark matter and baryonic processes most significantly influences galaxy evolution (Behroozi2019; Davies2019; Wang2025).
@@
Semi-analytical models (SAMs) offer a powerful framework for developing and calibrating such techniques. These models implement physically motivated prescriptions for galaxy formation within dark matter halo merger trees from N-body simulations (Prada2012; Somerville2015; Croton2016; Lagos2018; DeLucia2024). By producing realistic galaxy populations with known halo properties, SAMs allow us to assess the performance of different mass estimation techniques and quantify their associated uncertainties. Recent advances in SAMs, including improved treatments of gas cooling, star formation, and feedback processes, which has significantly enhanced their ability to reproduce observed galaxy properties across a wide range of environments (Klypin2016; Croton2016; Lacey2016; Lagos2018; Stevens2018; Henriques2020; Lagos2024; DeLucia2024).
@@
In this paper, we present two complementary approaches to improve observational halo mass estimates for galaxy groups. The first approach addresses limitations in traditional dynamical mass estimators via the virial theorem when applied to low-multiplicity groups by developing a corrective framework that accounts for systematic biases in groups with small velocity dispersions and projected radius measurements. The second method uses the correlation between baryonic mass and dark matter mass, probing the mass relationship between the three most massive galaxies in the halo and that of the halo mass.
@@
This paper is organised as follows. In Sections 2 & 3, we introduce the observational dataset in which we showcase applications of the halo mass relations and describe the SAMs used to calibrate/validate our relations. Section 4 introduces our two calibrated halo mass estimators; Section 4.1 presents the first approach utilising a modified virial theorem (MVT) to estimate the halo mass, and in Section 4.2 we present a summed stellar-halo mass relation (sSHMR) using the three most massive galaxies in a group as a proxy for halo mass. In Section 5, we apply the halo mass estimations to the observational data and demonstrate their performance and use cases. Finally, Section 6 summarises our findings and outlines future directions for halo mass estimation in upcoming large-scale surveys.
@@
2 Observational Data
The primary observational dataset used in this work is the Southern Galactic Pole (SGP) catalogue introduced in VanKempen2024. The SGP provides a highly complete spectroscopic sample of galaxies at $z<0.1$ across 376 deg2 ($340^{\circ}<\mathrm{RA}<26^{\circ}$, $35.3^{\circ}<\mathrm{Dec}<-25.8^{\circ}$), the Two-degree-Field Galaxy Redshift Survey (2dFGRS; Colless01) and Galaxy And Mass Assembly (GAMA; Driver2009) G23 survey regions. Redshifts are sourced primarily from 2dFGRS and GAMA, and are supplemented with the Six-degree Field Galaxy Survey (6dFGRS; Blake16), the 2-degree Field Lensing Survey (2dFLenS; Jones04), the 2MASS Redshift Survey (2MRS; Macri19), and the Million Quasars catalogue (MILLIQUAS; Flesch21) measurements. The SGP catalogue is comprised of 24,656 unique spectroscopic sources. These sources were cross-matched with photometry from the Wide-field Infrared Survey Explorer (WISE; Wright2010), combining both the WISE Extended Source Catalogue (WXSC; Jarrett13; Jarrett2019) and the point-source AllWISE catalogue (Cluver2014; Cluver2020). The matching between WISE photometry and that of the spectroscopic sources was approximately 93%, yielding mid-infrared measurements for 22,933 galaxies. The cross-matched WISE photometry enables robust estimates of stellar masses, derived from the W1-based relations of Jarrett2023, as well as star formation rates, calculated from W3 and W4 luminosities following the calibrations of Cluver2025.
As the SGP catalogue is constructed from multiple spectroscopic catalogues, the resulting dataset is heterogeneous in nature. To establish homogeneity for our analyses, we define three spectroscopic sub-samples: (1) SGP–2dF, comprising all galaxies with 2dFGRS photometry across the full SGP footprint; (2) G23–2dF, containing galaxies with 2dFGRS photometry restricted to the GAMA G23 region ($339^{\circ}<\mathrm{RA}<351^{\circ}$, $35^{\circ}<\mathrm{Dec}<-30^{\circ}$); and (3) G23–GAMA, consisting of galaxies with GAMA photometry within G23. Figure 1 presents an Right Ascension–Declination view of the homogeneous, WISE cross-matched sub-samples of the SGP dataset. The G23-GAMA sub-samples clearly demonstrate higher spectroscopic completeness, which is approximately 2.3 times greater than that of G23–2dF.
Figure 1: RA–Dec distribution of the three SGP sub-samples used for uniform completeness: SGP–2dF (red), G23–2dF (green), and G23–GAMA (orange). Comparing G23–2dF with G23–GAMA highlights the effect of spectroscopic completeness. Numbers in brackets denote sample sizes.
The SGP catalogue provides 1,413 galaxy groups, identified via a Python-based, graph-theory implementation of a friends-of-friends (FoF) algorithm, FoFpy (Lambert2020). FoFpy links galaxies when both their projected separations and line-of-sight velocity differences fall below scalable linking lengths, with a probabilistic cut used to reject links (see Lambert2020 for further details on the FoFpy algorithm). To ensure robust recovery of galaxy groups, a two-pass strategy was adopted: an initial FoF run with extended linking lengths captured larger groups, which were removed prior to a second pass with smaller linking lengths targeting smaller groups. The linking lengths were calibrated using mock lightcones constructed from the Millennium simulation (Springel05) and the Semi-Analytic Galaxy Evolution (SAGE) model (Croton2016). These mock lightcones were generated with 2dFGRS ($b_{J}=19.45$) and GAMA ($i=19.2$) magnitude limits. See VanKempen2024 for a full description of the construction of the SGP group catalogue.
3 Semi-Analytic Models of Galaxy Formation
@@
This section provides a comprehensive overview of the simulated datasets used in this study. We describe the suite of state-of-the-art SAMs of galaxy formation employed in this work. SAMs represent powerful theoretical tools for exploring the interplay between dark matter structure formation and baryonic physics. These models implement physically motivated prescriptions for key astrophysical processes within merger trees extracted from cosmological $N$-body simulations (see reviews by Baugh2006; Benson2010). By providing self-consistent galaxy populations with fully known dark matter halo properties, SAMs offer an ideal framework for developing and calibrating halo-mass estimation techniques. In this work, three independent SAMs are used with distinct roles: Shark (Lagos2018; Lagos2024) serves as the fiducial calibration model for the halo mass estimates, whereas SAGE and the GAlaxy Evolution and Assembly (GAEA; DeLucia2014; Hirschmann2016; DeLucia2024) models are used exclusively for validation and robustness testing. No parameters of the estimators are re-tuned on SAGE or GAEA. Each of these SAMs implement different physical prescriptions while operating on distinct $N$-body simulations. This multi-model approach enables assessment of robustness to variations in the underlying galaxy-formation physics and cosmology.
@@
To ensure comparability with the observations described in Section 2, all SAM outputs are post-processed into mock light cones with realistic sky coordinates and redshifts, including peculiar-velocity–induced redshift-space distortions. Apparent magnitudes are computed and an SGP-like selection is applied (e.g., GAMA G23 magnitude limit of $i<19.2$) to emulate the survey depth. Galaxy groups in the SAMs provides ground-truth memberships and halo masses while retaining observational selection effects. Estimator inputs are restricted to observables after the SGP-like selection, this methodology ensures that the calibration and validation of halo mass estimators are performed under conditions that closely mimic real, highly complete spectroscopic surveys, thereby enhancing the reliability and applicability of the methods to current and future observational datasets. As a caveat, a fainter magnitude limit of $Z<21.2$ is used for Shark in the development of the sSHMR (see Section 4.2.3).
3.1 SHARK